# MetaSym: A Symplectic Meta-learning Framework for Physical Intelligence

### v1

@@ 3.2 ActiveDecoder: Autoregressive Decoder
As mentioned in Section 1.1, we train an autoregressive decoder to model non conservative and realistic forces such friction, or air resistance, that break pure Hamiltonian symmetries. Unlike the SymplecticEncoder, which strictly enforces canonical updates, the ActiveDecoder can incorporate these extraneous phenomena. While the former is trained in isolation, we freeze its parameters during the decoder’s training.
We define $\mathbf{\hat{q}^{(i)}_{t}}$ and $\mathbf{\hat{p}^{(i)}_{t}}$ as the dissipative canonical coordinates, of a system $i$. An input consisting of $\{\mathbf{\tilde{q}^{(i)}_{t}},\mathbf{\tilde{p}^{(i)}_{t}},\Delta t,\mathbf{u^{(i)}_{t}}\}\in\mathbb{R}^{(2d+m+1)}$ is provided to a linear projection, where $\mathbf{u^{(i)}_{t}}\in\mathbb{R}^{m}$ represents the control input driving the system at time $t$. This linear layer which we call “ControlNet”, serves to map the input to a latent vector $\mathbf{z}_{d}\in\mathbb{R}^{2d}$ representing the non-conservative part of the system’s dynamics. We then apply a masked multi-head self-attention over the sequence $\{\mathbf{z}_{d}\}$ for autoregressive decoding. The masking allows us to model the causal dependencies of the decoder’s input sequence.
Next, we apply a cross-attention mechanism, augmented with a specialized meta-attention design. To achieve multi-system generalization and fast adaptation, we use a meta-learning setup akin to a bi-level optimization inspired by recent progress (Li et al., 2025). Specifically, we separate the ActiveDecoder’s parameters into: global parameters, $\theta_{AD}$, that remain shared across all systems, and local parameters, $\zeta_{i}$, that can be interpreted as system specific parameters. Specifically in the meta-attention:
@@
For each system $i$, we split the time-sequenced data, $\{\mathbf{\tilde{q}^{(i)}_{t}},\mathbf{\tilde{p}^{(i)}_{t}},\mathbf{u^{(i)}_{t}}\}$ into an adaptation set , $\mathcal{I}_{adapt}$, and a meta set , $\mathcal{I}_{meta}$. During the inner loop of the bi-level optimization, we hold $\theta_{AD}$ fixed and fine-tune $\zeta_{i}$ by performing a few steps of gradient descent. Each step iterates over $\mathcal{I}_{adapt}$ feeding the ActiveDecoder ($\Phi_{AD}$) with ground-truth $(\mathbf{q}_{t}^{(i)},\mathbf{p}_{t}^{(i)})$ and control $\mathbf{u}_{t}^{(i)}$, and minimizing an inner MSE loss:
$\mathcal{L}_{\mathrm{inner}}^{(i)}=\frac{1}{\mathcal{T}_{adapt}}\sum_{t\in\mathcal{I}_{\mathrm{adapt}}}\Bigl{\|}\Phi_{\text{AD}}$ $\bigl{(}\mathbf{q}_{t}^{(i)},\mathbf{p}_{t}^{(i)},\mathbf{u}_{t}^{(i)};\boldsymbol{\theta_{AD}},\boldsymbol{\zeta}_{i}\bigr{)}$ (6)
$-[\mathbf{q}_{t+1}^{(i)},\mathbf{p}_{t+1}^{(i)}]\Bigr{\|}^{2}.$
@@
$\mathcal{L}_{\mathrm{outer}}^{(i)}=\frac{1}{\mathcal{T}_{meta}}\sum_{t\in\mathcal{I}_{\mathrm{meta}}}\Bigl{\|}\Phi_{\text{AD}}$ $\bigl{(}\mathbf{q}_{t}^{(i)},\mathbf{p}_{t}^{(i)},\mathbf{u}_{t}^{(i)};\boldsymbol{\theta_{AD}},\boldsymbol{\zeta^{*}}_{i}\bigr{)}$ (7)
$-[\mathbf{q}_{t+1}^{(i)},\mathbf{p}_{t+1}^{(i)}]\Bigr{\|}^{2},$
propagates gradients only to $\theta_{AD}$. The overall training objective is then the sum over $\mathcal{L}_{\mathrm{outer}}^{(i)}$ for all systems $i$, over the batch.
We finalize the decoder with a Multi-Layer Perceptron (MLP) that outputs the position and momentum for the next time $\{\mathbf{\hat{q}^{(i)}_{t+1}},\mathbf{\hat{p}^{(i)}_{t+1}}\}$. The architecture is then used autoregressively for future predictions.
Figure 2: (Left) Represents the trajectory of the position and the momentum for arbitrary masses from the spring mesh system. The actual (orange line) and predicted (blue line) trajectories for the positions $q_{x}$ and $q_{y}$ along the $x$ and $y$ coordinates as well as the corresponding momenta ($p_{x},p_{z}$) are plotted. (Right) Plots the mean squared error of the trajectories for five random systems from the test set, along with each phase-space component’s median value.
@@ 3.3 Autoregressive Inference
In summary, by pairing our SymplecticEncoder (to ensure structure preservation) that provides a strong inductive bias with an Autoregressive Transformer-style Decoder (ActiveDecoder) equipped with meta-attention (to handle individual system variations) and end-to-end training with metalearning, our network, MetaSym, can rapidly adapt to diverse physical systems while guaranteeing physically consistent core dynamics. The following section presents a range of experimental results demonstrating superior performance compared to state-of-the-art approaches, especially in long-horizon stability and few-shot generalization to new system instances of varying complexity.
@@ 4 Results
In this section, we evaluate MetaSym’s performance, against a variety of State-of-The-Art (SoTA) models and benchmarks. In particular, we report the behavior of MetaSym against Transformers that have achieved impressive performance in modeling physical systems (Geneva & Zabaras, 2022) and DHNNs (Sosanya & Greydanus, 2022) that model systems via the Hemholtz decomposition, although they require gradient information and a numerical integrator to obtain trajectory data. Our model predicts the subsequent timesteps directly, which poses a much harder task that does not require numerical integration, which can be time-consuming and may hinder real-time performance. These results also reflect on the effectiveness of our design choices.
@@
As mentioned in Section 1.1, we choose to benchmark MetaSym for three extremely challenging and diverse datasets. Notably, we choose the spring mesh system, which tests scalability and closely resembles predicting the dynamics modeled using finite element methods. This dataset has provided the state-of-art testbed for large dimensional systems (Otness et al., 2021). We consider a novel dataset to demonstrate the robustness to noise and flexibility to model all types of physical systems with MetaSym by, considering an open quantum system undergoing heterodyne detection. Finally, we choose to predict the dynamics of a quadrotor, a long standing benchmark in robotics and challenging due to floating base dynamics (Bansal et al., 2016). The description of each experimental setup is laid in the next sections, however a more extensive description can be found in Appendix B.
4.1 Spring Mesh-System
For our first benchmark, we consider a 400 dimensional spring mesh system initially proposed by Otness et al. as part of the extensive benchmark for learning physical systems (Otness et al., 2021). The spring mesh system is provided as the most complex physical system among their benchmarks due to its high-dimensional state-space and dynamic couplings. This system is also used to model deformable volumes and the surfaces of complex objects that are vital to engineering tasks (Pfaff et al., 2021). In this setup, each spring position is discretized by nodes whose position is updated based on forces computed from spring extensions and compressions, thus subjecting the simulation to nontrivial communication and computation patterns between different node positions. For a more in-depth description of the benchmark, refer to Appendix B.1.
Figure 3: (Left) Represents the trajectory of the two measured quadratures obtained from heterodyne measurement for an arbitrary quantum system in the test set with measurement efficiency $\eta=0.86$ and measurement strength $\kappa$. The actual measurement trajectory (orange line) is plotted against the predicted measurement (blue line) by MetaSym. (Right) Plots the mean squared error of the trajectories along with the respective median value for five random systems with arbitrary initial states, and other randomized parameters from the test set.
@@ 4.2 Open Quantum System
To benchmark MetaSym on a novel open quantum dynamics dataset, we consider a parametric oscillator initialized in a coherent state (Gardiner & Zoller, 2010). The system Hamiltonian includes three contributions: the harmonic term, $\hat{H}_{\mathrm{osc}}=\omega\hat{a}^{\dagger}\hat{a}$, the squeezing term, $\hat{H}_{\mathrm{sqz}}=\frac{i\chi}{2}\left(\hat{a}^{\dagger 2}-\hat{a}^{2}\right)$, and the cubic driving term, $\hat{H}_{\text{cubic }}=\beta\left(\hat{a}^{3}+\hat{a}^{\dagger 3}\right)$. Here, $\hat{a}^{\dagger}$ and $\hat{a}$ represent creation and annihilation operators in a truncated Fock space of dimension $N$. Dissipation arises due to coupling with a thermal bath at rate $\gamma$ and average thermal occupation $\left\langle n_{\mathrm{th}}\right\rangle$. We assume that the system is continuously monitored via heterodyne detection, leading to a stochastic master equation for the conditional quantum state (refer to Appendix B.2). The primary observables are the real and imaginary components (quadratures) of the heterodyne current, denoted $X$ and $P$, respectively. For this benchmark, we generate the dataset by numerically solving the stochastic master equation for the setup described (Johansson et al., 2012). We produce a dataset with varying measurement efficiencies $\eta$, squeezing strength $\chi$, cubic term strength $\beta$, oscillator frequency $\omega$ and average thermal occupation $\langle n_{th}\rangle$ of the bath along with different initial states. This reflects real world experimental scenarios with variable $\eta$ due to sensors or non ideal measurement conditions. Refer to Appendix B.2 for a detailed description of the training setup and out-of-distribution inference setup that we used.
Figure 4: (Left) Represents the translational and angular phase-space evolution of the quadrotor, with training and test trajectories generated using the Crocoddyl trajectory optimization package (Mastalli et al., 2020). Each task is initialized with randomized initial conditions and a randomized terminal position over a 1.5s horizon. The ground-truth test trajectory (orange line) is compared to MetaSym’s predictions (blue line) indicating the good predictive capabilites of our model. (Right) Plots the average mean squared error of the trajectories for five randomly generated test-set systems with arbitrary initial conditions and terminal positions, along with randomized inertial system parameters.
@@ 4.3 Quadrotor
For our final benchmark, we consider a quadrotor system, a well-known robotic platform with various applications across different domains, since the agility and speed of such systems are challenging for current SoTA data-driven methods (Sanyal & Roy, 2023). The complexity of the systems originate from the floating-base dynamics and the discontinuous orientation representation (Zhou et al., 2019). The floating-base dynamics are prone to internal covariate shift during the training of data-driven models, since the magnitudes of the phase-space can vary across systems and time. Simultaneously, representing the quadrotor’s orientation with Euler angles introduces singularities, while unit-quaternions have a nonlinear constraint on their total magnitude which most times is enforced using penalties in the loss function. This can often lead to local-minima that can dramatically slow down or stop the training progress. Our physics-informed SymplecticEncoder, along with the meta-learning capabilities of the ActiveDecoder, presents increased robustness across these problems. To avoid additional terms in the loss function, we project the quaternions representing the floating-base’s orientation to their respective tangent-space. For a more elaborate description of our simulation setup, refer to Appendix B.3.

### v2

@@ 3.2 ActiveDecoder: Autoregressive Decoder
As mentioned in Section 1.1, we train an autoregressive decoder to model non conservative and realistic forces such friction, or air resistance, that break pure Hamiltonian symmetries. Unlike the SymplecticEncoder, which strictly enforces canonical updates, the ActiveDecoder can incorporate these extraneous phenomena.
We define $\mathbf{\tilde{q}^{(i)}_{t}}$ and $\mathbf{\tilde{p}^{(i)}_{t}}$ as the dissipative canonical coordinates, of a system $i$. An input sequence $\left(\mathbf{\tilde{q}^{(i)}_{t:c}},\mathbf{\tilde{p}^{(i)}_{t:c}},\Delta t,\mathbf{u^{(i)}_{t:c}}\right)\in\mathbb{R}^{(2d+m+1)\times c}$ with a context-window of c timesteps is provided to a linear projection, where $\mathbf{u^{(i)}_{t}}\in\mathbb{R}^{m}$ represents the m-dimensional control-input driving the system at time $t$. This linear layer, which we call “ControlNet”, serves to map the input to a latent vector $\mathbf{z}_{d}\in\mathbb{R}^{2d}$ representing the non-conservative part of the system’s dynamics. We then apply a masked multi-head self-attention over the sequence $\mathbf{z}_{d}$ for autoregressive decoding. The masking allows us to model the causal dependencies of the decoder’s input sequence.
Next, we apply a cross-attention mechanism, augmented with a specialized meta-attention design. To achieve multi-system generalization and fast adaptation, we use a meta-learning setup akin to a bi-level optimization inspired by recent progress [31]. We separate the ActiveDecoder’s parameters into: global parameters, $\theta_{AD}$, that remain shared across all systems, and local parameters, $\zeta_{i}$, that can be interpreted as system specific parameters. More specifically, in the meta-attention:
@@
For each system $i$, we split the time-sequenced data, $\mathbf{\tilde{q}^{(i)}},\mathbf{\tilde{p}^{(i)}},\mathbf{u^{(i)}}$ into an adaptation set , $\mathcal{I}_{adapt}$ containing $\mathcal{T}_{adapt}$ number of timesteps, and a meta set $\mathcal{I}_{meta}$ with $\mathcal{T}_{meta}$ number of timesteps. During the inner loop of the bi-level optimization, we hold $\theta_{AD}$ fixed and fine-tune $\zeta_{i}$ by performing a few gradient steps. Each step iterates over $\mathcal{I}_{adapt}$ feeding the ActiveDecoder ($\Phi_{AD}$) with ground-truth $(\mathbf{q}_{t:c}^{(i)},\mathbf{p}_{t:c}^{(i)})$ and control $\mathbf{u}_{t:c}^{(i)}$, and minimizing an inner MSE loss across $\mathcal{T}_{adapt}$ number of timesteps:
$\mathcal{L}_{\mathrm{inner}}^{(i)}=\frac{1}{\mathcal{T}_{adapt}}\sum_{t\in\mathcal{I}_{\mathrm{adapt}}}\Bigl{\|}\Phi_{\theta_{\text{AD}}}$ $\bigl{(}\mathbf{q}_{t:c}^{(i)},\mathbf{p}_{t:c}^{(i)},\mathbf{u}_{t:c}^{(i)};\boldsymbol{\theta_{AD}},\boldsymbol{\zeta}_{i}\bigr{)}-[\mathbf{q}_{t+1}^{(i)},\mathbf{p}_{t+1}^{(i)}]\Bigr{\|}^{2}.$ (4)
@@
$\mathcal{L}_{\mathrm{outer}}^{(i)}=\frac{1}{\mathcal{T}_{meta}}\sum_{t\in\mathcal{I}_{\mathrm{meta}}}\Bigl{\|}\Phi_{\theta_{\text{AD}}}$ $\bigl{(}\mathbf{q}_{t:c}^{(i)},\mathbf{p}_{t:c}^{(i)},\mathbf{u}_{t:c}^{(i)};\boldsymbol{\theta_{AD}},\boldsymbol{\zeta}^{*}_{i}\bigr{)}-[\mathbf{q}_{t+1}^{(i)},\mathbf{p}_{t+1}^{(i)}]\Bigr{\|}^{2},$ (5)
propagates gradients only to $\theta_{AD}$. The overall training objective is then the sum over $\mathcal{L}_{\mathrm{outer}}^{(i)}$ for all systems $i$, over the batch. We finalize the decoder with a Multi-Layer Perceptron (MLP) that outputs the position and momentum for the next timestep $\left(\mathbf{\hat{q}^{(i)}_{t+1}},\mathbf{\hat{p}^{(i)}_{t+1}}\right)$. The architecture is then used autoregressively for future predictions. The inherent high dimensionality of the ActiveDecoder’s internal representations, coupled with the lack of well-defined or tractable Hessian guarantees in this setting [32], motivates our adoption of this meta-learning paradigm over traditional methods such as MAML. By selectively adapting only a carefully chosen subset of ActiveDecoder’s parameters, we are able to minimize both the number of adaptation steps and the quantity of data required during the meta-learning process.
@@ 3.3 Autoregressive Inference
In summary, by pairing our SymplecticEncoder (to ensure structure preservation) that provides a strong inductive bias with an Autoregressive Transformer-style Decoder (ActiveDecoder) equipped with meta-attention (to handle individual system variations), MetaSym can rapidly adapt to diverse physical systems while guaranteeing physically consistent core dynamics. The proof of this strong inductive bias provided by the SymplecticEncoder and its effect on the training of the ActiveDecoder as part of MetaSym is provided by Appendix C. The sketch of this proof is as follows, the SymplecticEncoder $\Phi_{\theta_{\mathrm{SE}}}$ exactly preserves the form $d\Phi_{\theta_{\mathrm{SE}}}(\mathbf{x})^{\top}J\,d\Phi_{\theta_{\mathrm{SE}}}(\mathbf{x})=J$, while the ActiveDecoder can be written as $\Phi_{\theta_{\mathrm{AD}}}(\mathbf{z}_{c},\mathbf{u}_{t},\mathbf{d}_{t})$ where $\mathbf{u}_{t}$ and $\mathbf{d}_{t}$ represent the seperable, for the sake of this proof, control-input and dissipation. The Jacobian $\partial\Phi_{\theta_{\mathrm{AD}}}=I+A$ and $\|A\|\leq\rho$, where $\rho$ is shown empirically be bounded even under extreme deformation. Composing these maps gives $\Phi=\Phi_{\theta_{\mathrm{AD}}}\circ\Phi_{\theta_{\mathrm{SE}}}$ and hence $d\Phi=(I+A)\,d\Phi_{\theta_{\mathrm{SE}}}$. Expanding the symplectic form then yields, $d\Phi^{\top}J\,d\Phi=d\Phi_{\theta_{\mathrm{SE}}}^{\top}\bigl{(}J+A^{\top}J+JA+A^{\top}JA\bigr{)}\,d\Phi_{\theta_{\mathrm{SE}}}=J+O(\rho),$ which shows that $\Phi$ deviates from exact symplecticity by at most $O(\rho)$. An identical $O(\rho)$ argument on $\det d\Phi=\det(I+A)\,\det d\Phi_{\mathrm{SE}}$ establishes near–volume preservation, completing the proof that $\Phi$ is $O(\rho)$–near-symplectic.
@@ 4 Results
In this section, we evaluate MetaSym’s performance, against a variety of State-of-The-Art (SoTA) models and benchmarks, especially in long-horizon stability and few-shot generalization. In particular, we report the behavior of MetaSym against Transformers that have achieved impressive performance in modeling physical systems [19] and DHNNs [17] that model systems via the Hemholtz decomposition, although they require gradient information and a numerical integrator to obtain trajectory data. Our model predicts the subsequent timesteps directly, which poses a much harder task that does not require numerical integration, which can be time-consuming and may hinder real-time performance. These results also reflect on the effectiveness of our design choices.
@@
As mentioned in Section 1.1, we choose to benchmark MetaSym for three extremely challenging and diverse datasets. Notably, we choose the spring mesh system, which tests scalability and closely resembles finite element modeling of surfaces with different materials. This dataset has provided the state-of-art testbed for large dimensional systems [1]. Moreover, we utilize a well proven, reproducible dataset of an open-quantum system under undergoing heterodyne detection leading to representative quantum effects such as decoherence to demonstrate MetaSym’s robustness to noise and flexibility to model all types of physical systems. Finally, we choose to predict the dynamics of a quadrotor, a long-standing benchmark in robotics and challenging due to floating-base dynamics and sensor noise that is simulated using standard Gaussian additive noise [33]. The description of each experimental setup is laid in the next sections, however a more extensive description can be found in Appendix B. Finally, information regarding the utilized compute resources and training hyperparameters are provided in Appendices E & F.
Figure 2: (Left) Time evolution of the system’s position and momentum variables for a representative set of masses in the spring mesh. The orange curves represent the ground-truth trajectories for each phase-space coordinate type (i.e., $q_{x},q_{y},p_{x}$, and $p_{y}$), while the blue curves depict the model’s predictions. The close alignment between these trajectories underscores the model’s capacity to accurately capture the underlying long-term dynamics (600 timesteps) of the coupled spring system using a context window of 30 timesteps. (Right) Plots illustrating the mean squared error (MSE) of the time evolution of each phase-space coordinate type (dots) for five (A-E) spring-mesh systems in the test set. Each column encapsulates the spread of errors observed across all masses in the spring-mesh for a given phase-space coordinate across multiple timesteps, with the boxes marking their median values. The uniformly low median errors across all components demonstrate that the model generalizes effectively to different spring-mesh systems for all phase-space coordinates.
4.1 Spring-Mesh System
As our first benchmark, we use the 400-dimensional spring-mesh system introduced by Otness et al. [1], consisting of a $10\times 10$ lattice of point masses connected by elastic springs. Each of the 100 masses possesses two positional and two momentum degrees of freedom, yielding a 400-dimensional state space. This system, notable for its high dimensionality and complex dynamical couplings, serves as a canonical testbed for learning physical systems, including deformable volumes and surfaces relevant to engineering applications [34]. Node positions are updated via forces from spring extensions and compressions, resulting in nontrivial communication and computation patterns across the mesh. For a more in-depth description of the benchmark, refer to Appendix B.1. Unless stated otherwise, models are trained using teacher forcing for next-token prediction with a fixed context window of 30 timesteps. This choice aligns with prior configurations used in large-scale visual-language architectures (e.g., $\pi 0$: 50 timesteps [35], OpenVLA: 8 timesteps [36]). Notably, our models are significantly smaller, yet demonstrate strong capability for long-term, temporally consistent predictions within this horizon as the results in Figure 2 indicate.
Figure 3: (Left) Time evolution of the two quadratures measured via heterodyne detection for a representative quantum system in the test set, characterized by a measurement efficiency $\eta=0.86$ and measurement strength $\kappa$. The orange lines indicate the true measurement trajectories, while the blue lines display the model’s predictions (MetaSym). The close overlap between these trajectories highlights the model’s effectiveness in capturing the underlying quantum dynamics from heterodyne signals. The context window of the model is 30 timesteps. (Right) Plots showing the mean squared error (MSE) of the predicted quadratures (dots) for five randomly selected test systems. The consistently low median errors (boxes) across all tested systems underscore the robustness and generalization capabilities of the model.
@@ 4.2 Open Quantum System
To benchmark MetaSym on a novel open quantum dynamics task, we consider a parametric oscillator initialized in a coherent state [37]. The Hamiltonian includes a harmonic term $\hat{H}{{}_{\mathrm{osc}}}=\omega\hat{a}^{\dagger}\hat{a}$, a squeezing term $\hat{H}{{}_{\mathrm{sqz}}}=\frac{i\chi}{2}(\hat{a}^{\dagger 2}-\hat{a}^{2})$, and a cubic drive $\hat{H}{{}_{\text{cubic}}}=\beta(\hat{a}^{3}+\hat{a}^{\dagger 3})$, with $\hat{a},\hat{a}^{\dagger}$ the ladder operators in a truncated Fock space of dimension $N$. Dissipation arises via coupling to a thermal bath with rate $\gamma$ and mean occupation $\tilde{n}{{}_{\mathrm{th}}}$. Continuous heterodyne detection yields a stochastic master equation for the conditional state (Appendix B.2). The only experimentally accessible data are the real and imaginary quadratures $X$ and $P$ of the heterodyne current, extracted from single-shot trajectories, as the quantum state itself remains inaccessible. We simulate this setup by numerically solving the stochastic master equation [38], varying measurement efficiency $\eta$, squeezing strength $\chi$, cubic drive $\beta$, oscillator frequency $\omega$, bath occupation $\tilde{n}{{}_{\mathrm{th}}}$, and initial states. Training and out-of-distribution inference details are in Appendix B.2. As shown in Figure 3, MetaSym performs well in this strongly stochastic, non-classical regime.
Figure 4: (Left) Represents the translational and angular phase-space evolution of the quadrotor, with training and test trajectories generated using the Crocoddyl trajectory optimization package [39]. Each task is initialized with randomized initial conditions and a randomized terminal position over a 1.5s horizon (150 timesteps) with a context window of 30 timesteps. The ground-truth test trajectory (orange line) overlaps with MetaSym’s predictions (blue line) indicating the excellent predictive capabilites of our model. (Right) Plots summarizing the mean squared error (MSE) of the phase-space coordinates evolution for five randomly generated test systems (dots). The consistently low median errors (boxes) across all components underscore the robustness and generalization capabilities of the model.
@@ 4.3 Quadrotor
For our final benchmark, we evaluate on a quadrotor, a canonical yet challenging system due to its floating-base dynamics and discontinuous orientation representations [40, 41]. The floating base induces internal covariate shifts during training, as phase-space magnitudes could vary across time. Regarding its orientation representation, Euler angles suffer from singularities, while quaternions impose a unit-norm constraint that is typically handled via loss penalties, often leading to local minima and unstable training. To address this, we represent the pose of the quadrotor (i.e., position and orientation) using the tangent space of the $\mathbb{SE}(3)$ group [42], which avoids the explicit enforcement of the unit norm constraint. The results in Figure 4 verify that our framework effectively handles these challenges. Full simulation details are in Appendix B.3.